A 15kg box is being pulled with a force of 60N and is sliding with an acceleration of 4 m/s^2. Find the coefficient of kinetic friction. 


1
Expert's answer
2021-12-14T09:27:41-0500

"F=ma\to 60-\mu mg=ma\to \mu=(60-ma)/(mg)="


"=(60-15\cdot 4)/(15\cdot9.8)=0" . Answer